segment ab has endpoint A(-5,-4) and the midpoint M(2,0). What is the y-coordinate of point B?
Answer by stanbon(75887)

segment ab has endpoint A(-5,-4) and the midpoint M(2,0). What is the y-coordinate of point B?
--------
Let the coordinate form of B be (x,y).
Solve for x and for y::
(x-5)/2 = 2
(y-4)/2 = 0
-----
x = 4+5 = 9
y = 0+4 = 4
------
(x,y) = (9,4)
=================
Cheers,
Stan H.
-----------------